OpenStudy (nightmarenight):

Will fan & give medal for right answer????? Write a quadratic variation equation if g(x) varies directly with x2, and g(x) = 144 when x = 6.

OpenStudy (mertsj):

\[g(x)=kx^2\]

OpenStudy (mertsj):

Plug in the given values and solve for k

OpenStudy (nightmarenight):

does k=6?????? @Mertsj

OpenStudy (mertsj):

\[144=k(6)^2\] \[144=36k\] \[4=k\]